# Hyperradiance, photon blockade and concurrence in a pair of qubits inside a driven cavity

### v1

@@ A Generation of hyperradiance by a two-photon drive in the absence of Kerr medium
$\begin{split}R&=\left(\frac{-8\eta^{2}(4g^{2}+16d^{2}+(\kappa+\gamma)^{2})}{|A+iB|^{2}}+\frac{4\eta^{2}(|C|^{2}+8g^{2}|(2\Delta-i\gamma)|^{2})}{|A^{\prime}+iB^{\prime}|^{2}}\right)\times\\ &\frac{D}{8\eta^{2}(4g^{2}+16d^{2}+(\kappa+\gamma)^{2})}\end{split}$ (12)
where the coefficients $A,A^{\prime},B,B^{\prime},C,D,E$ are mentioned in Appendix. A in Eq. (22). Figures. 5(a) and Fig. 5(b) show that the master equation and perturbative analytical result of Eq. (12) match quite well.
To examine whether the enhanced hyperradiance has anything to do with two-qubit entanglement we plot concurrence between qubits as a function of $\eta/\gamma$ and $\Delta/\gamma$ for $\chi=0$ in Fig.4 (b). Fig.4 (a) shows that at $\Delta/\gamma=0,\pm\frac{\sqrt{6}g}{2}$ the two qubits are highly entangled. It can be understood from Fig. 5(c) where two-qubit concurrence $C(\rho_{q})$ is plotted as a function of $\Delta/\gamma$ for $\eta=0.1\gamma$. The concurrence goes down when the system is not in the hyperradiant regime or in subradiant regime. Therefore, it is worth noting that the radiance and concurrence has a one-to-one connection as far as the low driving regime is concerned. In Fig. 4(c) we plot the minimum of the squeezing-parameter Min$[S_{\theta}]$ as a function of $\Delta/\gamma$ and $\eta/\gamma$. There are three significant peaks at $\Delta=0,\pm\sqrt{6}g/2\gamma$ for weak pumping. The corresponding squeezing angle $\theta_{s}$ is plotted in Fig. 4 (d). The rest of the parameters in Fig. 4 are $g_{1}=10\gamma,g_{2}=10\gamma,\kappa=0.5\gamma$ chosen to fulfill the strong coupling CQED limit.
We display the Wigner functions of the cavity field in Figs. 6(a-c) for three typical parameter sets chosen from the Fig. 4 (b) to verify squeezing of the field. In contrast to the wider range of fluctuations as Fig.6(b) shows, the Wigner functions of Figs. 6 (a) and Fig. 6(c) show elliptical profile in the parameter regime of squeezing, showing that the distribution is elongated in one particular direction. Figures. 6 (a) and 6(c) exhibit squeezing along almost vertical and horizontal directions, respectively, depending on the various pump field detunings.
To illustrate the nonclassical photon number distribution in Fock space, we plot the corresponding Klyshko’s criteria distribution $K_{n}$ in Figs. 6(d–f). The odd-even oscillation of $K_{n}$ alternates up and down mirroring the expected even-odd photon distribution of squeezed states. It is important to note that unlike 6 (b) that corresponds to the resonant case i.e $\Delta=\pm\sqrt{6}g/2\gamma$, the off-resonant cases in Fig. 6(a) and Fig.6(c) show a rapid decay of the even-odd oscillations with the increase of photon number. It means that the squeezed states generated in the off resonant cases involve mainly two-photon excitations where in resonant cases higher order even number transition processes may take place.
In order to understand the $n$-photon blockade, $\log_{10}g^{(n)}(0)$ (n=2,3) is plotted in the plane of $\Delta/\gamma$ and $\eta/\gamma$ in Figs.7 (a) - (b) which show that none of the correlations goes below unity ($\log_{10}g^{(n)}(0)<0$). So it is impossible to find a single-photon or two photon blockade in this entire parameter range. The reason for the absence of single photon blockade can be explained from the dressed-state picture. The transition from $\psi^{(0)}\rightarrow\psi_{0,\pm}^{(1)}$ is not allowed as they are not coupled by two photon transitions. The value of two and three photon correlations are very large near $\eta=0$. Needless to say in these positions, the average photon numbers are very low making the statistical measurements a difficult task.
So far, we have studied the squeezing and hyperradiance caused by two-photon and multiphoton processes. But, no single or two photon blockades have been found. We next introduce Kerr nonlinear medium in the cavity to see if we can achieve blockades and how the squeezing and hyperradiance are modified due to the nonlinearity.
B The effect of inra-cavity Kerr-nonlinearity: Photon blockade
Now we introduce Kerr-nonlinearity in order to explore nonlinearity induced photon blockade [31, 52]. In Fig.8 (a) $R$ is plotted as a function of $\eta/\gamma$ and $\Delta/\gamma$. The three peaks obtained correspond to the three transitions in the dressed-state levels. To confirm this behaviour, we have plotted $R$ in Fig. 9 as a function of $\Delta/\gamma$. Both perturbative and master equation based numerical values of $R$ are in good agreement. Moreover, the incorporation of nonlinearity $\chi$ does not affect the radiance behaviour much. In Fig.8 (b) Min$[S_{\theta}]$ and (c) $\theta_{s}$, the minimum squeezing angle is plotted in the plane of $\eta/\gamma$ and $\Delta/\gamma$ with $\chi=6$. Unlike Fig. 4, this figure shows three peaks at different detunings. Because the presence of nonlinear medium has modified the dressed state structure. The peaks appeared in Fig.8 (a), (b) and (c) are at $\frac{E_{0}^{(2)}}{2}$, $\frac{E_{2}^{(2)}}{2}$ and $\frac{E_{3}^{(2)}}{2}$ respectively. It is because the transitions $\psi_{0}^{(0)}\rightarrow\ \psi_{0}^{(2)},\psi_{2}^{(2)},\psi_{3}^{(2)}$ are allowed and it can be confirmed by calculating the transition dipole matrix elements between the two dressed states. Likewise, In Fig.8 (b) the significant cocurrence is obtained at $\frac{E_{0}^{(2)}}{2}$, $\frac{E_{2}^{(2)}}{2}$ and $\frac{E_{3}^{(2)}}{2}$, the same positions where hyperradiance is obtained. It can be again confirmed by Fig. 9 (c) as discussed in previous section. This justifies the fact that in detunings corresponding to the resonant transitions between the dressed states, the entanglement between the qubits is significant, making the collective radiation from the two qubits enhanced. It is to be noted that the Min[$S_{\theta}$] value has almost the same minimum as it is obtained for $\chi=0$. The minimum squeezing angle $\theta_{s}$ is usually dependent on the detunings and not on the driving strength as before. Rest of the parameters are same as in Fig. 4.
Now, to explore the $n$-photon blockade, we plot a phase space curve in Fig. 10 (b) and Fig. 10 (c). In Fig. 10 (b), the yellow region describes the condition when both $g^{(2)}(0)$ and $g^{(3)}(0)$ are less than unity. In this region, we also confirm that $g^{(3)}(0)<g^{(2)}(0)$. In Fig. 10 (c) the three yellow regions depict the possibility of two photon blockade characterised by $g^{(3)}(0)<1,g^{(2)}(0)>1$. In the blue region of (a) and (b), the condition concerning the yellow region is not satisfied. Also, we have ensured that in the yellow region, $\bar{n}$ is appreciable, otherwise the detection of correlation will be a difficult work.
To ensure the existence of $n$-photon blockade, we choose the most suitable parameters from 10 (b) and (c) and explored the variation as a function of detuning $\Delta/\gamma$. In Fig. 11 (a) and (b) $\bar{n}$ (in the left y axis) and $\log_{10}g^{(n)}(0)$ (in the right y axis) vs. $\Delta/\gamma$ is plotted. In Fig. 11 (a) a strong two photon blockade regime $\Delta/\gamma=\frac{E_{0}^{(2)}}{2}$ and for $\eta=0.63\gamma$ characterised by $g^{(2)}(0)>1$ and $g^{(3)}(0)<1$ is found. In this regime $g^{(2)}(0)\approx 1$ and $g^{(3)}(0)\approx 0.01$ which indicates an enhanced two-photon blockade. A moderately strong two photon blockade is also found at $\Delta/\gamma=\frac{E_{2}^{(2)}}{2}(g^{(3)}(0)\approx 0.43),\frac{E_{3}^{(2)}}{2}(g^{(3)}(0)\approx 0.1)$ and for $\eta=0.63\gamma$. In Fig. 11 (b) we show single photon blockade characterised by $g^{(3)}(0)<g^{(2)}(0)<1$ is observed at $\Delta/\gamma=\frac{E_{0}^{(2)}}{2}$ and for $\eta=1.53\gamma$.
The foregoing results can also be confirmed by comparing the photon-number distributions and the Poisson distributions. In order to verify (a), the two-photon blockade, we plot $\frac{P_{n}-{\Pi}_{n}}{\Pi_{n}}$ with respect to photon number in Fig. 11 (c) and found out that only two photon probabilities are enhanced ($P_{2}>\Pi_{2}$) and all the higher order photon probabilities are suppressed ($P_{2}>\Pi_{2}$ for $n>2$). To verify single photon blockade in (b) we plot the same as (c) in 11(d) and found out that only single photon probabilities are enhanced ($P_{1}>\Pi_{1}$) and all the higher order photon probabilities are suppressed ($P_{1}>\Pi_{1}$ for $n>1$). It is to be noted that in both the cases (a) and (b), we plotted $\bar{n}$ to confirm the presence of appreciable amount of photon number at the respective points for the sake of experimental realization. The rest of the parameters are same as in 10.
Now, to corroborate our results, we have judiciously chosen a parameter regime and explored all the statistical properties together. In Fig. 13 (a) and (b), we have plotted $R$ and $C(\rho_{q})$ as a function of $\Delta/\gamma$ and $\eta/\gamma$. We show that at $\Delta/\gamma=\frac{E_{0}}{2}$ the highest peak appears in the $R$ vs. $\Delta/\gamma$ curve has a finite concurrence. We also investigate squeezing in the same regime. The Wigner distribution shows eliptical profile in the X-Y space in the diagonal direction that indicates the emitted radiation is squeezed. Squeezing is indeed verified by the Kleysko’s parameter $K_{n}$ which shows prominent even- odd oscillations in the photon number space. Lastly, we plot $g^{(n)}(0)$ and $\bar{n}$ as a function of $\Delta/\gamma$. $\bar{n}$ shows three distinct peaks at $\frac{E_{0}}{2},\frac{E_{2}}{2},\frac{E_{3}}{2}$. In these positions $g^{3}(0)<1$ and $g^{2}(0)>1$ indicating a 2P blockade. In these three positions the $g^{2}(0)\approx 2,23.5,3$ and the $g^{(3)}(0)$ is achieved as low as $g^{(3)}(0)\approx 0.05,0.6,0.02$. So, $g^{(2)}(0)$ is two order of magnitude higher than $g^{(3)}(0)$ at $\frac{E_{0}}{2}$ which indicates an enhanced 2P blockade. The strong 2P blockade is again confirmed by plotting $\frac{P-\mathcal{P}}{\mathcal{P}}$ as a function of photon number. This quantity shows a peak at $n=2$ and all other photons to be negative indicating a two photon emission exploiting the possibility of higher number of photons being present in the output radiation. Thus our system can act as a hyperradiant two-photon squeezed light generator with a pair of entangled qubits.
Figure 1: Schematic diagram of a pair of qubits inside a cavity that may contain a Kerr nonlinear medium. Here, we have shown a pair of two-level atom as qubits. The cavity is driven by a two-photon drive with strength $\eta$ and frequency $\omega_{d}$. Atoms have the transition frequency $\omega_{a}$ and are coupled to the cavity with strength $g_{i}$. $\kappa$ and $\gamma$ are the decay rates of cavity and the atom, respectively.
Figure 2: Dressed-state level diagram for the coupled two-qubit-cavity system without Kerr nonlinearity. The Rabi frequency of the pump field is labelled as $\eta$. Only few photon transitions are shown in this figure.
Figure 3: Dressed-state energies are plotted as a function of $g$ with $\chi=6\gamma$ (solid line) and $\chi=10\gamma$ (dashed line).
Figure 4: (a) The radiance witness $R$, (b) concurrence $C(\rho_{q})$ and (c) Min[$S_{\theta}$] are plotted as a function of $\eta/\gamma$ and $\Delta/\gamma$. The greener area corresponds to the squeezing of the field. The rest of the parameters are $g_{1}=10\gamma,g_{2}=10\gamma,\kappa=0.5\gamma$, $\chi=0$. (d) Squeezing angle $\theta_{s}$ is plotted on the plane of the $\Delta/\gamma$ and $\eta/\gamma$. The blank areas in (b) corresponds to zero concurrence. The blank areas in figures (c) and (d) indicate that the minimum squeezing parameter is non-zero.
Figure 5: (a) $R$ and (b) $\langle a^{\dagger}a_{i}^{M(C)}\rangle$ is plotted as a function of $\Delta/\gamma$ by solving both Eq. (12) (solid) and the master equation 3 (dashed). Here $M(C)$ represents master equation and complex hamiltonian where $i$ stands for single qubit and two qubits in the cavity. (c) Concurrence as a measure of entanglement between two qubits are plotted as a function of $\Delta/\gamma$. For (a), (b) and (c) the parameters are $\eta/\gamma=0.1$ and rest of the parameters are same as in fig. 4.
Figure 6: The density plots of Wigner functions (a), (b) and (c) and the corresponding Klyshko’s parameter ($K_{n}$) (d), (e) and (f) are depicted for three sets of $\eta/\gamma,\Delta/\gamma$ values as mentioned at the top of the each sub-figures (a), (b) and (c) chosen from Fig. 4(b). The other parameters are same as in Fig. 4
Figure 7: The second and third order HBT correlation functions $g^{(2)}(0)$ and $g^{(3)}(0)$ are plotted as a function of $\Delta/\gamma$ and $\eta/\gamma$. The other parameters are same as in Fig. 4. As $g^{(2)}(0)$ and $g^{(3)}(0)$ are not less than 1, one- or two- photon blockades do not exist.
Figure 8: (a) $R$, (b) $C(\rho_{q})$ (c) Min[$S_{\theta}$] is plotted as a function of $\eta/\gamma$ and $\Delta/\gamma$. The greener area indicates the squeezing nature of the field. (d) squeezing angle $\theta_{s}$ is plotted on the plane of the normalized detuning $\Delta/\gamma$ and driving strength $\eta/\gamma$ for the same coupling of two-qubit-cavity system. The white area of figure (b) shows $C(\rho_{q})=0$. The blank areas in both the figure (c) and (d) indicates that the minimum squeezing parameter min[$S_{\theta}>0$]. Here, the parameters are $g_{1}=10\gamma,g_{2}=10\gamma,\kappa=0.5\gamma$, $\chi=6\gamma$.
Figure 9: (a) $R$ and (b) $\langle a^{\dagger}a_{i}^{M(C)}\rangle$ is plotted as a function of $\Delta/\gamma$ by solving both Eq. (12) (solid) and the master equation 3 (dashed). Here $M(C)$ represents master equation and complex hamiltonian where $i$ stands for single qubit and two qubits in the cavity. (c) Concurrence as a measure of entanglement between two qubits are plotted as a function of $\Delta/\gamma$. For (a),(b) and (c) the parameters are $\eta/\gamma=0.1$ and rest of the parameters are same as in fig. 8.
Figure 10: (a) $\bar{n}$ , (b) the single-photon and (c) two-photon blockade regime(yellow portion) are shown on the plane of the normalized detuning $\Delta/\gamma$ and driving strength $\eta/\gamma$ for the same coupling with $\chi=6\gamma$. The other parameters are same as in fig. 8.
Figure 11: (a) The average photon number (blue dashed curve), (b) the field correlation functions $g^{(2)}(0)$ (solid) and $g^{(3)}(0)$ (dotted), are plotted as a function of $\Delta/\gamma$, for $\eta=0.63\gamma$ and (b) $\eta=1.53\gamma$. The other parameters are same as in fig. 8. (c) and (d) depict deviations of the photon-number distribution from the standard Poisson distribution when $\Delta/\gamma=\frac{E_{0}^{(2)}}{2},\frac{E_{2}^{(2)}}{2}and\frac{E_{3}^{(2)}}{2}$. (a) and(c) show two-photon blockade and (b) and(d) show single-photon blockade at detuning $\frac{E_{0}^{(2)}}{2}$.
Figure 12: (a) $R$ and (b) $C(\rho_{q})$ are plotted as a function of $\Delta/\gamma$. (c) Wigner function is plotted in $X$ and $Y$ space, (d) Kleyshko’s parameter $K_{n}$ is plotted as a function of $n$. (e) $\bar{n}$ (solid), $g^{(n)}(0)$ ($n=2(\rm{dashed}),3(\rm{dashed-dotted})$) vs. $\Delta/\gamma$ and (f) $\frac{P-\mathcal{P}}{\mathcal{P}}$ vs. $n$ are shown. For all these figures $\eta=0.4\gamma$. For (c), (d) and (f) the detuning is kept constant at $\frac{E_{0}}{2}$. The rest of the parameters are same as in Fig. 8.
IV Conclusions
In summary, we have studied the quantum properties of a CQED system with a pair of qubits inside a cavity driven by a two-photon drive. We have shown that the qubits give rise to hyperradiance which can be linked with the two-qubit entanglement in low driving regime. We have further demonstrated that the emitted hyperradiant light can be squeezed in quadrature. To confirm squeezing, we have calculated Wigner distribution of the field and Kleyshko’s parameter. The Wigner distribution shows eliptical profile while the Kleyshko’s parameter exhibits even-odd oscillations as a clear signature of squeezing. We have also explored the influence of an intra-cavity Kerr-nonlinear medium on the system. The nonlinear medium induces photon blockade that can be employed to control the photon statistics of the field. We have shown that the hyperradiant field can be generated with single as well as two-photon blockades in suitable parameter regimes. The hyperradiant field with two-photon blockade is found to be quadrature-squeezed. Interestingly in these regimes, the system posseses a finite two-qubit concurrence confirming the entanglement between the qubits. Our findings may have potential applications in quantum communication and quantum networking.
acknowledgement
AD is thankful to Indian Association for the Cultivation of Science, Kolkata, India and DST, Govt. of India for research fellowship and Subhanka Mal for fruitful scientific discussions.
Appendix A Analysis using non-hermitian Hamiltonian approach
In what follows, we present approximate analytical expressions for the radiance $R$ of the transmitted field. Under the scenario of weak driving ($\eta<<g,\kappa$) the total excitation of the qubit-cavity system can be assumed to be two as in Refs. [53, 34]. In this circumstance, the joint wave function of the pair of qubits and photons system can be reasonably approximated in the two-photon manifold with the ansatz
$\begin{split}|\psi_{2}(t)\rangle&=C_{gg,0}(t)|gg,0\rangle+C_{gg,2}(t)|gg,2\rangle+C_{gg,1}(t)|gg,1\rangle+C_{+,1}(t)|+,1\rangle+\\ &C_{+,0}(t)|+,0\rangle+C_{ee,0}(t)|ee,0\rangle\end{split}$ (13)
Because the radiance requires at least a single qubit present in the cavity, we treat single qubit-cavity system in the two photon space also with the ansatz wave function,
Under the weak driving limit, the qubit-cavity system remains predominantly in the ground state. So, the terms arising as $2a\rho a^{\dagger}$ and $2\sigma_{j}\rho\sigma_{j}^{\dagger}$ in the Lindblad master equation can safely be ignored [54, 55]. This leads to the non-Hermitian Hamiltonian given by
and
For the sake of simplicity, we take $g_{1}=g_{2}=g$, $\Delta_{a}=\Delta_{c}=\Delta$ and $\gamma_{1}=\gamma_{2}=\gamma$. Based on the time-dependent Schrödinger equation $i\frac{\partial|\psi\rangle}{\partial t}=H^{i}_{nh}|\psi\rangle$, the dynamics of the above probability amplitude coefficients $C_{mn}$ incorporating $H^{1}_{nh}$ and $\psi_{1}(t)$can be described by a set of coupled equations as follows
$i\hbar\dot{C_{g,1}}$ $=$ $\Delta C_{g,1}+gC_{e,0}-i\frac{\kappa}{2}C^{1}_{g,1}$
$i\hbar\dot{C_{e,0}}$ $=$ $\Delta C_{e,0}+gC_{g,1}-i{\gamma}C_{e,0}$
$i\hbar\dot{C_{e,1}}$ $=$ $2\Delta C_{e,1}+\sqrt{2}gC{g,2}-i(\frac{\kappa}{2}+\frac{\gamma}{2})C^{1}_{e,1}$ (17)
. The coefficient $C_{mn}$ stands for the probability amplitude of the corresponding state $|m,n\rangle$. In the weak driving limit, we have the relationship $C_{g,0}>>C_{g,2},C_{e,1}>>C_{g,1},C_{e,0}$ since it is a two photon drive. So, we assume that $C_{g,0}\approx 1$ [35]. Again using Eqs. 16 and 13, we get the coupled differential equations of the form
The coefficient $C_{lmn}$ stands for the probability amplitude of the corresponding state $|lm,n\rangle$. Again, considering weak driving we assume that $C_{gg,0}\approx 1$. We solve these dynamical equations in the steady state. For the sake of simplicity, we have defined two effective detuning term as $\Delta_{k}=\Delta-\frac{i\kappa}{2}$ and $\Delta_{l}=\Delta-\frac{i\gamma}{2}$ which incorporate damping rates in the system. The solution of Eq.(17) and Eq.(18) are given by,
The mean photon number for single qubit-cavity system and with pair of qubit in steady state is given by,
$\langle{\hat{a}}^{\dagger}a\rangle_{1,i}$ $=$ $|C_{g,2}|^{2}+2|C_{e,1}|^{2}$
$\langle{\hat{a}}^{\dagger}a\rangle_{2}$ $=$ $|C_{gg,1}|^{2}+2|C_{gg,2}|^{2}$ (21)
The detailed analytical expression of $R$ is thus obtained as given in Eq. (12)
@@ Appendix A Analysis using non-hermitian Hamiltonian approach
$E$ $=$ $(16\Delta^{2}\kappa^{2}+8g^{2}\kappa\gamma+2\kappa^{3}\gamma+\kappa^{2}\gamma^{2})$ (22)
To verify the above analysis, we compare the numerical result by solving the master equation with the result form Eq. (12) for both $\chi=0,6$. As shown in Fig., these two results matches well except the amplitude or intensity. This is because the states included in the complex hamiltonian approach excludes the gain pathways that is incorporated in the density matrix master equation approach via the Lindblad term. Thus the numerical results has a little high magnitude than the analytical one.
The equal-time second-order correlation function can be expressed as,
@@
It has a complicated expression. We have plotted both the numerically and analytically obtained $g^{(2)}(0)$ Both the results are in good agreement except the amplitudes do not match exactly because of the same reason mentioned earlier.

### v3

@@ B Radiance and Concurrence
$\begin{split}R&=\left(\frac{-8\eta^{2}(4g^{2}+16\Delta^{2}+(\kappa+\gamma)^{2})}{|A+iB|^{2}}+\frac{4\eta^{2}(|C|^{2}+8g^{2}|(2\Delta-i\gamma)|^{2})}{|A^{\prime}+iB^{\prime}|^{2}}\right)\times\\ &\frac{D}{8\eta^{2}(4g^{2}+16\Delta^{2}+(\kappa+\gamma)^{2})}\end{split}$ (7)
@@
$E$ $=$ $(16\Delta^{2}\kappa^{2}+8g^{2}\kappa\gamma+2\kappa^{3}\gamma+\kappa^{2}\gamma^{2})$
To explore qubit-qubit entanglement we make use of the concurrence [54] defined by
C Criteria for $n$ photon blockade ($n$PB)
[…]
$P_{n}\geq\Pi_{n},P_{m>n}<\Pi_{m>n}$ (10)
where $\Pi_{n}$ is the Poisson distribution given by $\Pi_{n}=\langle n\rangle^{n}\exp(-\langle n\rangle)/n!$. $\langle n\rangle$ is the average number of photons given by $\langle n\rangle=\rm{Tr}[a^{\dagger}a\rho]$. According to Eq. (10), the $n$PB effect pertains to an increase of the likelihood of $n$ photons while suppressing the probabilities of the photon numbers greater than $n$. The other criterion is based on $g^{(n)}(0)$, an equal-time nth-order correlation function. The $n$PB effect must meet the following requirements
$g^{(n)}(0)\geq 1,g^{(n+1)}(0)<1$ (11)
meaning that the field has $(n+1)$th order sub-Poissonian photon statistics and the $n$th order poissonian or super-Poissonian photon statistics. In particular, the 2PB effect satisfies the correlation functions $g^{(2)}(0)\geq 1$ and $g^{(3)}(0)<1$. The 1PB effect can be characterized by $g^{(2)}(0)<<1$. We have calculated $g^{(2)}(0)$ using the numerical solution of the master Eq. 3 as well as the perturbative analytical solution as obtained in Appendix- A through complex hamiltonian approach. Following the later procedure the equal-time second-order correlation function can be expressed as
@@ C Criteria for $n$ photon blockade ($n$PB)
where $C_{gg,2}$ and $C_{+,1}$ are the probability amplitude of the states $|gg,2\rangle$ and $|+,1\rangle$ as defined in Appendix A.
D Nonclassicality of the field
In order to seek possible connection between $n$-photon blockade and the nonclassical nature of the cavity field, we calculate three quantities, namely squeezing parameter $S_{theta}$, Klyshko’s number $K_{n}$ and radiance witness $R$.
where $X_{\theta}=(ae^{-i\theta}+a^{\dagger}e^{i\theta})$ is the linear combination of the hermitian quadrature operators $X=\frac{(a+a^{\dagger})}{\sqrt{2}}$ and $Y=\frac{i(a^{\dagger}-a)}{\sqrt{2}}$. Any negative value of $S_{\theta}$ implies squeezing nature of the field characterized by narrowing of the quadrature distribution compared to that of a coherent state. To verify the nonclassicality of the photon number distribution, we use the Klyshko’s criterion [48]
$K_{n}=\frac{(n+1)P_{n-1}P_{n+1}}{nP_{n}^{2}}$ (14)
E Two-qubit dressed-state picture
In order to explain radiance-concurrence-photon blockade connection, it may be instructive to work in dressed-state picture. In the absence of coherent drive and decay processes, we develop the dressed-state structure for the system in order to have some insight of the physical processes taking place in the system. Throughout our calculations, we have assumed that qubit-cavity coupling for both the atoms are same $g_{1}=g_{2}=g$. We have defined the collective two-qubit basis states as $|gg\rangle$,|±$|\pm\rangle=\frac{|ge>\pm|eg\rangle}{\sqrt{2}}$ and $|ee\rangle$ to describe the dynamics in the dressed-state picture. In the single-photon space the eigenvalues and eigenvectors (dressed states) are obtained as $|\psi_{0}^{(1)}\rangle=-|+,0\rangle$ with $E_{0}^{(1)}=0$ and $|\psi^{(1)}_{\pm}\rangle=\frac{|gg,1\rangle\pm|+,0\rangle}{\sqrt{2}}$ with $E_{\pm}^{(1)}=\pm\sqrt{2}g$. In the two-photon space, the eigenvectors are $|\psi^{(2)}_{0}\rangle=-\frac{1}{\sqrt{3}}|gg,2\rangle+\frac{\sqrt{2}}{\sqrt{3}}|ee,0\rangle$ and $|\phi^{(2)}_{0}\rangle=-|+,1\rangle$ with corresponding eigenvalue $E_{0}^{(2)}=0$. The another eigenvector is $\psi_{\pm}^{(2)}=\frac{1}{\sqrt{3}}|gg,2\rangle\pm\frac{1}{\sqrt{2}}|+,1\rangle+\frac{1}{\sqrt{6}}|ee,0\rangle$ with corresponding eigenvalue $E_{\pm}^{(2)}=\sqrt{6}g$. Due to the symmetric coupling ($g_{1}=g_{2}=g$) the antisymmetric Dicke state $|-,n\rangle$ is decoupled in the $n$photon space ($n=1,2$). In order to know the allowed transitions between the dressed states in the presence of a two-photon dive, we calculate the transition dipole matrix elements between the dressed states as shown in Fig. 2. In the one-photon space, Kerr nonlinearity ($\chi$) will not affect the dressed states. It will act on the states with photon numbers $n\geq 2$. So, the one-photon dressed states and dressed energies in the presence of Kerr nonlinearity will be same as those in the absence of the nonlinearity. In the two-photon space the hamiltonian in the basis $|gg,2\rangle,|\pm,1\rangle$ and $|ee,0\rangle$ is given by,
. The dressed energies and dressed states are obtained by diagonalizing this Hamiltonian. We have plotted the dressed energies of the system for $\Delta=0,\chi=6\gamma,10\gamma$ in Fig. 3. The dressed energies for $\Delta=0$ with $\chi=6\gamma,g=10\gamma$ are $E_{0}^{(2)}=-21.35,E_{1}^{(2)}=0,E_{2}^{(2)}=3.80,E_{3}^{(2)}=29.55$ and the corresponding dressed states are
$|\psi_{1}^{(2)}\rangle$ $=$ $-|+,1\rangle$
Figure 2: Dressed-state level diagram for the coupled two-qubit-cavity system without Kerr nonlinearity. Three most significant transitions are shown in this figure. All these three transitions have the same matrix element $\sqrt{\frac{2}{3}}\eta$ with $\eta$ being the pump field Rabi frequency.
Figure 3: Dressed-state energies are plotted as a function of $g$ for $\Delta=0$ with $\chi=6\gamma$ (solid line) and $\chi=10\gamma$ (dashed line).
III Results and discussions
We first present the results in the absence of Kerr medium in subsection III(A) and then discuss the effects of Kerr nonlinearity ($\chi$) in subsection III(B). All the frequency quantities are scaled by the qubit decay constant $\gamma$.
Here we explore radiance, concurrence between two qubits, squeezing and nonclassicality of the field. Our results suggest a plausible connection between the radiance which is the property of the field and the concurrence that is a measure of entanglement between the qubits. However, there exists no photon blockade in the absence of any Kerr nonlinearity.
The radiance witness $R$ is plotted as a function of the scaled detuning $\Delta/\gamma$ and driving strength $\eta/\gamma$ in Fig. 4 (a). Near $\Delta=0,\pm\sqrt{6}g/2\gamma$ the system exhibits hyperradiant behaviour due to the multiphoton excitations to the dressed eigenstates as illustrated in the dressed-state level diagram of Fig. 2. The strengths of the transition $|\psi^{(0)}\rangle\rightarrow|\psi_{0,\pm}^{(2)}\rangle$ are same and it is $\sqrt{\frac{2}{3}}\eta$. So, unlike [25, 31] where only $|\psi^{(0)}\rangle\rightarrow|\psi_{0}^{(2)}\rangle$ transition was allowed, we have three distinct hyperradiance peaks at three different detunings. In fact, we find regimes where $R\approx 50$ which is substantially greater than the previously reported results so far [28, 26, 27, 25, 29]. To examine whether the enhanced hyperradiance has anything to do with two-qubit entanglement, we plot concurrence between the qubits as a function of $\eta/\gamma$ and $\Delta/\gamma$ for $\chi=0$ in Fig.4 (b). It shows that at $\Delta/\gamma=0,\pm\frac{\sqrt{6}g}{2}$ the two qubits are highly entangled. Now, to verify the nonclassicality, we plot the minimum of the squeezing-parameter Min$[S_{\theta}]$ as a function of $\Delta/\gamma$ and $\eta/\gamma$ in Fig. 4(c). There are three significant peaks at $\Delta=0,\pm\sqrt{6}g/2\gamma$ for weak pumping. The corresponding squeezing angle $\theta_{s}$ is plotted in Fig. 4 (d). The rest of the parameters in Fig. 4 are $g_{1}=10\gamma,g_{2}=10\gamma,\kappa=0.5\gamma$ chosen to fulfill the strong coupling CQED limit. Figures. 5(a) and 5(b) show $R$ and $\langle a^{\dagger}a\rangle$ as a function of $\Delta/\gamma$ at $\eta=0.4$ using both perturbative analytical result given by Eq. (7) and numerically solving master Eq. (3). As shown in these figures, these two results match quite well except slight mismatch in the amplitude or intensity. This is because the states included in the complex hamiltonian approach excludes the gain pathways unlike the master equation approach of Eq. 3. Figure 4(b) can be understood from Fig. 5(c) where two-qubit concurrence $C(\rho_{q})$ is plotted as a function of $\Delta/\gamma$ for $\eta=0.1\gamma$. Three peaks appearing at the transition frequencies as shown in Fig. 5(a) correspond to moderate concurrence. The concurrence goes down when the system is not in the hyperradiant regime or in subradiant regime. Therefore, it is worth noting that the radiance and concurrence seems to have a one-to-one connection as far as the low driving regime is concerned.
Refer to captionRefer to caption
Refer to captionRefer to caption
Figure 4: (a) The radiance witness $R$, (b) concurrence $C(\rho_{q})$ and (c) Min[$S_{\theta}$] are plotted as a function of dimensionless pump strength $\eta/\gamma$ and dimensionless detuning $\Delta/\gamma$. (d) Squeezing angle $\theta_{s}$ is plotted on the plane of the $\Delta/\gamma$ and $\eta/\gamma$. The rest of the parameters are $g_{1}=10\gamma,g_{2}=10\gamma,\kappa=0.5\gamma$ and $\chi=0$. The blank areas in (b) corresponds to zero concurrence.
Refer to captionRefer to captionRefer to caption